A Bibliography of Publications about the MINIX Operating System

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

06 February 2023 Version 2.29

Title word cross-reference

1/4in [Tan91b]. 1000 [Lin09]. 11 [EG17]. 11th [JE06]. 13th [Ano90c, Ano90a]. 1988 [IEE88b]. 1990 [Win91].

2 [Tan91b]. 2002 [Fra02]. 2006 [ACM06a]. 20th [IEE94]. 22nd [Ano89a]. 28th [Ano95].

3 [Ahm08, HBG^{+}06d, HBG^{+}06e, Her10, Lin09, Meu06, Swi10, Sze11, TAB^{+}10, Tan15b, Tan15c, dS08, vM07, vdK09].

4in [Tan91b].

512K [Tan87d].

640K [Tan87e]. 68000 [Mei91]. 68000-rechner [Mei91].

’87 [Ano87]. ’88 [IEE88a].

Accountability [YC05]. ACM [ACM88, ACM06a, Fra02]. ACSAC [JE06]. Ada [NKN93]. Addition [Ang91, Lun17b]. Address [GKT12]. Advances [JE06]. Aerospace [IEE88b]. Afraid [HBG^{+}06a]. aid [AEG^{+}91]. Alternative [Ano90b, Yag90, vMAT14, GLG93]. AMD [Chr96, Nol04]. AMD-K5 [Chr96]. Amsterdam [Ahm08]. Anais [Sil88]. Analysis [Fra02, ACM06b, Gre90]. Andrew [Hof10, Sev14]. Annual [ACM88, Ano89a, Ano95]. application [ABFL92]. Applications [IEE88b, Ior13, IEE88a, VOJ^{+}92]. Applying [FPA06]. approach [DSX06, Ola97, dJKH93]. Approximations [Nan88b]. April [Ano89b, Ano95]. Architecture [HBG^{+}06b, Alt06b, Alt06a, Chr96, IEE94, JE06, NCCN88, Wil98]. Architectures [Pri12]. Artificial [IEE88a]. Asia [JE06]. Asia-Pacific [JE06]. ATARI [TSM88, Dur89, GD89b, Tan91c]. Atlanta [ACM88]. authorization [FPA06, FBM88]. Automated [GT12, GCT13]. Automatic [GKT13b, Giu14, Ior13]. Autumn [Ano92]. av [Aas89, Vik93].

Back [GIKT13]. Bad [vMAT13a]. balance [TCJ94]. Based [AvMT10b, App14, McG97, vMAT13b, Ang91, AvMT12, Kac89, PN92, Wil98]. Battling [vMAT13a]. Bay [IEE88a]. Belgium [Ano89b]. Benchmarking [GKT13a]. besturingssysteem [Lun17a]. binaries [Tan87d, Tan87e]. Bits [vMAT13a]. Block [AvMT10a, McG97]. Block-Based [McG97]. Block-level [AvMT10a]. Brasileira [Sil88]. Brazilian [Sil88]. brings [Her05b]. broadcast [Kac89]. Brussels [Ano89b]. build [Ano90d]. Building [App14, HBT06, Her10, Meu06]. Byte [Hof10].

C [Ahm08, NKN93]. Cache [AvMT14, vMAT13a]. Caches [AvMT14]. Caching [AvMST13]. call [Her90]. Can [THB06]. Canada [ACM06a, ACM06b]. Carolina [Kan92, Win91]. Causeway [IEE88a]. challenges [Ano89b]. check [FPA06]. Checkpointing [VGBT13]. Checksums [vMAT13a]. China [JE06]. Chip [Lin11, bta17]. Clone [Tan87b, Tan87c, Tan87f]. CLONIX [Ano90d]. Cloud [Lin11]. Code [Tan87b, Tan87c, Tan87f]. Collection [Ahm08]. College [Ano89a, Ano95]. communication [Cus88, TC91].

Comparing [Ahm08]. Compiler [Ahm08]. Components [HBT14]. Computacao [Sil88]. Computer [ACM88, Ano90c, Ano90a, DW08, IEE88a, IEE88b, Lin11, MHY^{+}95, DSX06, DTW07, JE06]. computers [Ano90b, Yag90]. Computing [Ano89a, Sil88, VOJ^{+}92, Ano95]. concerns [Lun17b]. Conference [ACM88, Ano87, Ano90c, Ano90a, IEE88a, IEE88b, IEE94, ACM06a, Ano89b, JE06, Ano92]. Congres [Ano89b]. Congress [Sil88]. Congresso [Sil88]. considerations [Mag88]. Construction [HBG^{+}06c]. Cooperative [GT09]. cost [Ano90b, Yag90]. Countering [HBG^{+}08]. Course [Hay89]. courses [AEG^{+}91, Tan87f]. CPU [bta17]. CQUAL [FPA06]. Crash [vMAT12, vMAT13b]. Crashed [GCT10]. Creating [Alt06b, Alt06a, Nan88b]. CT [Fer91]. CT-MiniFrame [Fer91]. Current [TAB^{+}10].

database [MR90, Ren90]. DC [Ano90c]. Dead [Her06, HBG^{+}06a, AvMT10a]. Dealing [HvMA^{+}09]. December [IEE88a, IEE88b]. Dependability [GKT13a, vMAT11]. Dependable [AvMT10b, GT09, GKT13a, HBG^{+}06c, Her10, HVBT12, MFH^{+}09, YC05]. Design [CAH90, Ger06a, Ger06b, Mag88, Ram88, Tan87a, Tan88c, TW97, ACM06a, Ola97, TW06, TW09]. desk [Gre90]. Despite [HBT06]. Developing [Chr96]. Development [RT93, Chr96, Lar90, Ram88, TT93]. Device [HBT06, Her06, HBG^{+}06a, HBG^{+}07a, HBG^{+}09, OK95, SABL04, Her05b, KPG93, Nol04]. didactical [AEG^{+}91]. differences [NKN93]. different [Gre90]. Disabling [EG17]. discretionary [FBM88]. Disk [Wei92, dJKH93]. Distortion [vdKGT14a]. Distributed [McG97, YC05, GH89, Her89, Mag88, MR90, MM91, Nan88b, Nan88a, Ram88, Ren90, San90, TCJ94].

Distribution [OW02]. DMINIX [TC91]. do [Sil88]. Down [AvMT14]. Driver [HvMA^{+}09, OK95, Nol04, Sze11]. Drivers [HBT06, Her06, HBG^{+}06a, HBG^{+}07a, HBG^{+}09, SABL04, Her05b]. Dutch [Lun17a]. Dynamic [Vee09].

East [Ano92]. EDFI [GKT13a]. Education [DW08, DSX06, DTW07]. Effectiveness [ABFL92]. Efficient [GKT12, Pri12, VGBT13, vHvAvMT11]. electronic [Hof10]. Embedded [Hei05]. End [vMAT11]. End-to-End [vMAT11]. Engine [bta17, Tys17]. England [IEE94]. Enhanced [GKT12]. Enhancement [Guh89]. enhancements [Her05b]. Enhancing [MFH^{+}09]. Environment [RT93, TT93]. EUROMICRO [IEE94]. European [Ano89b]. EurOpen [Ano92]. EurOpen. [Ano92]. EUUG [Ano87, Ano89b]. Evaluating [vdKGT14a]. evaluation [GLG93]. Everywhere [AvMT14]. Evolutionary [MM91]. example [DTC90]. Excelsior [IEE88a]. Exclusivity [AvMT14]. Execution [RT93, TT93]. Experience [Har90]. Experiencing [AEG^{+}91]. Experiments [GKT13a, vdKGT14a, vdKGT14b]. Extending [KPG93]. extension [FBM88]. eyebrows [Lun17b].

facility [TCJ94]. Failure [HBG^{+}07a, HBG^{+}07b, Vee09]. Failure-Resilient [HBG^{+}07b]. Failures [HvMA^{+}09, vdKGT14b]. Falls [Ano95]. Fast [HVBT12]. Faster [HBT13a, HBT13b]. Fault [GIKT13, GKT13a, HBG^{+}06b, HBG^{+}09, Her10, vdKGT14a, vdKGT14b]. Fault-Injection [vdKGT14a]. Fault-Resilient [HBG^{+}06b]. Fault-tolerant [GIKT13]. Faults [OW02]. February [ACM88]. File [AvMT10b, AvMT11, AvMST13, App14, Ger06a, McG97, Vee09, vMAT13b, GH89, Her89, Hd90, San90, Wei92, dJKH93, Ger06b]. File-Based [AvMT10b, App14]. File-Level [AvMST13]. Finding [dS08]. Fine [GKT12]. Fine-grained [GKT12]. firmware [Fra19]. First [Win91]. Flash [AvMT12, AvMST13]. Flash-based [AvMT12]. Flexibility [App14]. Flexible [AvMT11, vMAT14, FBM88]. Florida [IEE88b]. flow [FPA06]. flow-sensitive [FPA06]. Flushing [AvMT14]. Flying [Chr96]. focus [ACM88]. Fourth [IEE88b]. Future [GIKT13, Ano90c, Ano90a].

Generic [OK95]. Georgia [ACM88]. GNU [Ahm08]. Goes [Ano92]. grained [GKT12].

hardware [GD89a]. help [Vik93].

Heterogeneous [HBT13a, HBT13b, Pri12]. Highly [HBG^{+}06c, HBG^{+}06d]. Hits [AvMT14]. hjelp [Vik93]. Hong [IEE88a]. Host [AvMST13]. Host-Side [AvMST13]. hot [dS08]. Hotel [Ano90c, IEE88a]. HP [Aas89]. HP-Minix [Aas89]. HP-Minux [Aas89]. Hybrid [AvMT14].

IBM [Tan87d, Tan87e, Tan87b, Tan87c, Tan88b, Tan88a, Tan91b]. Impact [Sev14, Tan15a]. Implement [Her90]. Implementation [Chi95, Cus88, Fer91, Ger06a, GLG93, Nan88a, San90, Tan87a, Tan88c, TW97, Tiw90, ACM06a, lC95, CAH90, Fre90, Ger06b, Kob89, Lou92, MM91, Ola97, TW06, TW09, Xu95]. Implementing [Lin09, Lin11, Wai95]. Improving [Lak04, dJKH93]. In-memory [VGBT13]. independency [Alt06b, Alt06a]. independent [Chr96]. Industrial [OW02]. Information [Ano90a, Ano90c]. Injection [GKT13a, vdKGT14a, vdKGT14b]. instruction [Koc90]. Instructional [DW08, DSX06, DTW07]. instruments [Chr96]. Integrated [vMAT11, vMAT12, Ola97]. Integrating [AvMT12]. integration [IEE94]. Intel [EG17, Lin09, Lun17b, Tan17, Tys17, bta17]. intelligence [IEE88a]. intense [MHY^{+}95]. interaction [Ash97, MHY^{+}95]. interface [LG88]. International [Fra02, IEE88a, Kan92, Win91]. Interprocess [TC91]. introduction [Byf10]. Invariants [GCT13]. Investigating [vdKGT14b]. IOMMU [Sze11]. IP [Tiw90]. IPC [HBG^{+}08]. Isolation [HBG^{+}09]. ISSTA [Fra02]. Italy [Fra02].

jisso [TWC98]. July [Fra02]. June [ACM06a, Kan92, Win91]. just [Gre90].

K5 [Chr96]. Keep [HVBT12]. kernel [Cus88, Her05b]. Key [Ano90a, Ano90c]. Kit [Ahm08]. know [Lun17b]. Known [Ano17]. Kong [IEE88a].

lab [Har90]. Laboratories [DW08, AAS94, DTW07]. language [ACM06a, NKN93]. languages [ACM06b]. Large [OW02]. Last [Hof10]. learned [Tan16]. learning [Ano90d]. Least [Ano17]. Lessons [Tan16]. Letter [Tan17]. Level [AvMST13, AvMT10a]. Lightweight [HBT06, vMAT14]. Linda [CG93, Vik93]. Lindex [Vik93]. Linux [Wil98]. Live [GT09, GIKT13, GKT13b, Giu14, Ior13]. Liverpool [IEE94]. Load [TCJ94]. Logical [RT93, TT93, dJKH93]. Loris [AvMT10b, AvMT11, AvMST13, App14, vMAT11, vMAT12, vMAT13a, vHvAvMT11]. low [Ano90b, Yag90]. low-cost [Ano90b, Yag90].

M3P [NCCN88]. M3P-project [NCCN88]. Machine [RT93, TT93]. Macintosh [Gre90]. main [Lun17b]. Make [THB06]. Management [Tys17, bta17, vHvAvMT11, GD89a, KPG93, Lak04, Ren90]. manager [Nan88a]. manual [TSM88, TKS92]. Masses [Gre90]. ME [EG17]. measure [Ang91]. Measurement [Meu06]. mechanism [KK88, Lou92]. mechanisms [FBM88]. meets [CG93]. MegaST [Tan91c]. Memory [GD89a, Lak04, VGBT13]. message [Ang91, Ash97, Kob89]. Metadata [vHvAvMT11]. Method [HBT06]. methodology [Wil98]. Microkernel [Her05b]. Microkernels [Hei05, Hof10]. Microscope [Ano90b, Yag90]. microwaves [MHY^{+}95]. might [Lun17b]. migration [Lou92]. MiniFrame [Fer91]. minikkusu [TWC98]. Minix [Her90, Mei91, Ahm08, Alt06b, Alt06a, Ang91, Ano90d, ABFL92, Byf10, lC95, CAH90, Cus88, DTC90, FPA06, Fre90, GH89, Ger06a, Ger06b, GLG93, GD89b, Har90, Her05b, HBG^{+}06d, HBG^{+}06e, Her10, Her89, Hof10, Kac89, Kel06, Kob89, Koc90, Lak04, Lar90, LG88, Li93, Lin09, Lun17b, Lun17a, Mag88, MR90, Meu06, Nan88b, Nan88a, OK95, Ram88, Ros88, San90, Sev14, Smi91, Swi10, Sze11, Tan87b, Tan87c, TSM88, Tan88b, Tan91a, Tan91c, Tan91b, TAB^{+}10, Tan15a, Tan15b, Tan15c, Tan16, TCJ94, Tys17, Vai96, Vik93, Wai95, Xu95, Yan95, bta17, dS08, vM07, vdK09, Aas89, AEG^{+}91, Ano90b, AAS94, Chi95, CG93, DSX06, Fer91, Guh89, Hay89, How02, KPG93, Lin11, Lou92, Nol04, Ola97, Tan87d, Tan87e, Tan88a, TKS92, Tiw90, Vee09]. Minix [Wil98, Yag90]. MINIX/THL [Koc90]. MinixPPC [Alt06b, Alt06a]. MINNET [Kac89]. Minux [Aas89]. Mode [Swi10, EG17]. Model [GT09, Alt06b, Alt06a, Her90, Ros88]. Modern [Tan01]. Modular [AvMT10b, AvMT11, HBG^{+}06e, vHvAvMT11]. Modularity [App14]. Modules [vMAT13b]. Monitor [RT93, TT93]. Monitoring [GCT13]. Most [Ano17, Lun17b, Lun17a]. moving [Her05b]. multi [Dur89, Hd90]. multi-transputer [Hd90]. multi-user [Dur89]. multicast [Cus88, TC91]. Multicore [Pri12]. Multicores [HBT13a, HBT13b]. Multilevel [AvMT14]. Multimedia [vM07]. multiprocessor [PN92, Vai96]. Multis [Dur89]. Multiserver [HBG^{+}08, HBT14, MFH^{+}09, Pri12]. Multitasking [Gre90].

Namespace [vMAT13b]. National [Ano90c, Ano90a]. Need [Hei05]. Net [HVBT12]. network [Kac89]. Networking [HVBT12]. networks [Ash97]. ni [TWC98]. nonlinear [MHY^{+}95]. North [Kan92, Win91]. Norwegian [Aas89, Vik93]. novel [DSX06]. NRDNIX [Ren90].

October [Ano90c]. Omni [Ano90c]. On-chip [bta17]. Ontario [ACM06a, ACM06b]. Open [Fra19, Tan17]. Operating [Ano90b, GKT12, GKT13b, GD89b, Hay89, Her05a, Her05b, HBG^{+}06b, HBG^{+}06c, HBT06, HBG^{+}06d, HBG^{+}07b, HBG^{+}08, Her10, How02, Kui12, MFH^{+}09, Meu06, OK95, Ola97, RT93, Tan87a, Tan88c, TW97, THB06, TW06, TW09, Tan15c, Yag90, Aas89, AEG^{+}91, Ang91, Ano90d, AFL91, ABFL92, AAS94, CAH90, Cus88, DSX06, Fre90, Guh89, Har90, Kob89, Koc90, Lun17b, Lun17a, Mag88, MM91, Nan88b, PN92, Ram88, Tan87f, Tan01, Tiw90, TC91, TT93, Yan95]. operativsystem [Aas89]. Operetingu [TWC98]. Optimal [HBT14]. ordinary [Ano90b, Yag90]. oriented [Kob89]. Orlando [IEE88b]. Ottawa [ACM06a, ACM06b]. oyobi [TWC98].

Pacific [JE06]. packages [Dur89]. Page [vMAT13a]. pain [GD89a]. Palais [Ano89b]. panacea [GD89a]. paper [San90]. Parallel [VOJ^{+}92]. Park [Kan92, Win91]. Parsing [Cox01]. Part [Ano90b, Yag90]. passing [Ang91, Ash97, Kob89]. path [Kan92, Win91]. PC [Gre90, Tan87b, Tan87d, Tan87e, Tan87c, Tan88b, Tan88a, Tan91b]. PC-AT [Tan87d, Tan88a]. PCnet32 [Nol04]. Peachtree [ACM88]. Performance [Ahm08, Meu06]. personal [Ano90b, Yag90]. placement [FPA06]. plasmas [MHY^{+}95]. Platform [Alt06a, Alt06b, Vai96]. Plaza [ACM88]. PLDI [ACM06a]. policy [Ros88]. Polymorphic [Kui12]. populairste [Lun17a]. popular [Lun17b, Lun17a]. Port [Alt06a, GD89b, Alt06b]. portering [Aas89]. Portierung [Mei91]. Porting [Aas89, Kel06, Li93, Vai96, vdK09]. PowerPC [Alt06b, Alt06a]. Practical [GCT13]. Pro [Lin09]. Pro/1000 [Lin09]. procedure [Her90]. Proceedings [ACM06a, Ano89a, Ano90a, Ano92, Ano89b, Fra02, IEE88a, IEE94, ACM88, ACM06b, Ano87, Ano90c, JE06]. Process [vMAT12, vMAT13b, Ang91, GLG93, Kob89, Lou92, Wil98]. process-based [Ang91]. processes [Xu95]. processing [Smi91]. processor [Chr96]. Program [GCT13]. Programming [ACM06a, ACM06b, HBG^{+}06e, Alt06b, Alt06a, NKN93]. project [Lar90, NCCN88]. projects [How02]. Protection [MFH^{+}09, FBM88]. prototype [Win91, Kan92]. Prototyping [Kan92, Win91, AFL91, ABFL92]. PS [Tan91b]. PS/2 [Tan91b]. put [Gre90].

Q&A [Hof10]. QEMU [vdK09]. quality [Her05b].

RAID [AvMT10a]. raising [Lun17b]. Randomization [GKT12]. Rapid [Kan92, Win91, AFL91]. Real [Smi91, KK88, Wai95]. Real-time [Smi91, Wai95]. rechner [Mei91]. Recovering [SABL04]. Recovery [vMAT12, vMAT13b]. reduces [Her05b]. reference [TKS92]. reimplementasjon [Vik93]. reimplementation [Vik93]. Reincarnation [Her06]. Reliability [HBG^{+}06f]. Reliable [GT12, Her05a, HBT06, HBG^{+}06d, HBT13b, Pri12, THB06, Tan15c]. remote [Her90]. Reorganizing [HBG^{+}06f]. Repairing [HBG^{+}06d]. Report [TAB^{+}10]. Research [Her05a, Kan92, TAB^{+}10, Win91]. Reservations [MFH^{+}09]. Resilience [HBG^{+}07a]. Resilient [HBG^{+}06b, HBG^{+}07b]. Resistance [Vee09]. Resource [MFH^{+}09, FBM88, Nan88a]. revision [Her05b]. revisited [How02]. riron [TWC98]. Roadmap [HBG^{+}07b]. Role [YC05]. Roma [Fra02]. RS232 [Kac89]. running [Lun17b]. Runs [bta17, Ano90b, Tys17, Yag90].

S [Hof10, Sev14]. Safe [GT12, GKT13b, Giu14, Ior13]. SCCS [Ano89a]. schedulers [GLG93]. Scheduling [HBT14, Swi10, KK88]. Science [ACM88, IEE88a]. SD [Ano95]. Second [Kan92]. Secure [GT12, Hei05, THB06, Tan15c]. Security [Ano90a, DW08, GKT12, IEE88b, ACM06b, Ano90c, DSX06, Ros88, DTW07]. SEED [DTW07, DW08]. sekkei [TWC98]. Self [HBG^{+}06d]. Self-Repairing [HBG^{+}06d]. sensitive [FPA06]. September [IEE94, JE06]. Server [Vee09, Her89, Hd90]. services [Wai95]. Shanghai [JE06]. sharing [FBM88]. shisutemu [TWC98]. Shoreham [Ano90c]. shortening [Kan92, Win91]. sic [Alt06b]. Side [AvMST13]. SIGPLAN [ACM06a]. SIGSOFT [Fra02]. Silence [vdKGT14b]. Silent [vdKGT14b]. simple [Her90]. simulation [MHY^{+}95]. Single [Lin11]. Sink [AvMT14]. Sioux [Ano95]. Sixteenth [ACM88]. size [Her05b]. Slower [HBT13a, HBT13b]. Small [Ano89a, Ano95]. Sociedade [Sil88]. Society [Sil88]. sockets [Chi95]. Software [Fra02, OW02, vdK09, ACM88]. Solaris [Wil98]. Soundness [vdKGT14b]. Source [Tan87b, Tan87c, Fra19, Tan87f]. sources [Tan87d, Tan87e]. Space [GKT12, Her05b, MHY^{+}95]. SPARC [Wil98]. specification [Kan92, Win91]. spots [dS08]. Spring [Ano87, Ano89b]. SSDs [AvMT12]. ST [Dur89, GD89b, TSM88, Tan91c]. Stack [AvMT10b, AvMT12, App14, HvMA^{+}09, HVBT12, vMAT11, vMAT12]. Standards [Ano90a, Ano90c]. State [GT12, GIKT13, GCT13]. Status [TAB^{+}10]. steps [MM91]. Storage [AvMT10b, AvMT12, App14, HvMA^{+}09, vMAT11, vMAT12]. strongly [Her05b]. structure [LG88]. structures [Wei92]. Study [Xu95, Yan95, KK88]. Suite [DW08, DTW07]. Summary [Her05a]. SunOS [AAS94, Chi95]. Support [vM07, FBM88, TC91]. Supporting [RT93, TT93]. swapper [CAH90]. swapping [Fre90, Kob89]. Symposium [Ano89a, Fra02, Ano95]. System [Ger06a, GKT12, GD89b, Her05b, HBG^{+}06b, HBG^{+}06c, HBG^{+}06d, HBG^{+}06e, HBG^{+}07b, Her10, HBT14, Kan92, MFH^{+}09, Meu06, OK95, OW02, Tan88c, Tan15c, Win91, vMAT12, vMAT13b, Aas89, Ang91, ABFL92, AAS94, lC95, CAH90, Cus88, DSX06, Fre90, GH89, Ger06b, Guh89, Hd90, Kob89, Koc90, KK88, Lun17b, Lun17a, Mag88, MR90, MM91, NKN93, Nan88b, PN92, Ram88, Ren90, San90, Smi91, Tiw90, TC91, TCJ94, Yan95, IEE94]. Systematic [Hof10]. Systems [Ano90b, Ano90a, GKT13b, Hay89, Hei05, Her05a, HBT06, HBG^{+}08, HBT13b, Kui12, McG97, Pri12, RT93, Tan87a, TW97, THB06, Yag90, YC05, AEG^{+}91, Ahm08, Ano90c, Ano90d, AFL91, Har90, Hof10, How02, JE06, Koc90, Ola97, Tan87f, Tan01, TW06, TW09, TT93, Wei92, dJKH93].

talks [Hof10]. Tanenbaum [Hof10, Sev14]. TCP [Tiw90]. TCP/IP [Tiw90]. Techniques [VGBT13]. template [AFL91, ABFL92]. Temporal [MFH^{+}09]. ter [Lun17a]. Testing [Fra02]. thanks [Lun17b]. theory [IEE88a, MM91]. Thinking [Hof10]. THL [Koc90]. Threats [HBG^{+}08]. Time [GIKT13, KK88, Smi91, Wai95]. Time-traveling [GIKT13]. Tolerance [Her10]. tolerant [GIKT13]. Tool [GKT13a, AAS94]. Tools [Meu06, Gre90]. training [Koc90]. Transaction [vMAT13b]. Transaction-Based [vMAT13b]. Transfer [GT12, GIKT13]. transformation [Mag88]. Transputer [VOJ^{+}92, Hd90, PN92]. transputer-based [PN92]. traveling [GIKT13]. Triangle [Kan92, Win91]. TRIX [PN92]. True [Her05b]. Trusted [DTC90].

undocumented [EG17]. UNIX [Ano92, Ano89b, HBG^{+}06f, Tan87b, Tan87c, Tan87f, Ano90b, FBM88, Hd90, Ior13, Wei92, Yag90]. Unreliable [HBT06]. Update [GT09, GT12, GIKT13, GKT13b, Giu14, Ior13]. Updates [Vee09]. USA [Kan92, Win91]. Use [Pri12, AAS94]. Used [Ano17]. User [Swi10, Dur89, Her05b, LG88, Wil98]. user-interface [LG88]. user-space [Her05b]. Using [Ash97, GCT13, Hay89, vdKGT14b, DSX06, Kac89, Lou92, Ola97].

ved [Vik93]. verify [FPA06]. VFS [Ger06b]. via [EG17, MFH^{+}09]. VIII [Sil88]. Virtual [Ger06a, Ger06b]. Virtualization [AvMT11, vMAT14, vdK09]. visualization [Ash97]. VM [Li93]. Volume [AvMT11]. voting [Hof10]. Vulnerability [GCT13].

Washington [Ano90c]. wereld [Lun17a]. Westin [ACM88]. Who [HBG^{+}06a]. Widely [Ano17]. Window [lC95]. without [Chr96]. worked [DTC90]. Working [HVBT12]. Workshop [Kan92, Win91, ACM06b]. World [Ano17, Lun17b, Lun17a].

x86 [Ahm08, Chr96]. Xen [Kel06]. XML [Cox01]. XT [Tan88b, Tan91b].

years [Tan16]. yoru [TWC98].

References

Aas:1989:HMP

[Aas89] Gisle Aas. HP-Minix: portering av et operativsystem. (norwegian). [HP-Minux: Porting of an operating system]. Hovedoppgave i datafag (computer science thesis), Universitetet i Tromsø, Tromsø, Norway, 1989. 161 pp.

Ashton:1994:SMT

[AAS94] P. Ashton, D. Ayers, and P. Smith. SunOS Minix: a tool for use in operating system laboratories. Australian Computer Science Communications, 16(1):259–269, 1994. CODEN ACSCDD. ISSN 0157-3055.

Archer:1992:EOS

[ABFL92] M. Archer, J. Bock, D. Frincke, and K. Levitt. Effectiveness of operating system prototyping from a template: application to MINIX. In Kanapoulos [Kan92], pages 55–66. ISBN 0-8186-3040-X. LCCN QA76.76.D47 I598 1991. IEEE Catalog No. 92TH0454-9.

ACM:1988:PFS

[ACM88] ACM, editor. Proceedings, focus on software / 1988 ACM Sixteenth Annual Computer Science Conference, February 23–25, the Westin, Peachtree Plaza, Atlanta, Georgia. ACM Press, New York, NY 10036, USA, 1988. ISBN 0-89791-260-8. LCCN QA 76.758 A26 1988.

ACM:2006:PPA

[ACM06a] ACM, editor. PLDI 2006:

Proceedings of the 2006 ACM SIGPLAN conference on Programming language design and implementation 2006, Ottawa, Ontario, Canada, June 11–14, 2006. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-320-4. LCCN ???? ACM order number 548060.

ACM:2006:PWP

[ACM06b] ACM, editor. Proceedings of the 2006 workshop on Programming languages and analysis for security, Ottawa, Ontario, Canada. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-374-3. LCCN ????

Aguirre:1991:EMD

[AEG^{+}91] G. Aguirre, M. Errecalde, R. Guerrero, C. Kavka, G. Leguizamon, M. Printista, and R. Gallard. Experiencing Minix as a didactical aid for operating systems courses. Operating Systems Review, 25 (3):32–39, July 1991. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Archer:1991:TRP

[AFL91] M. Archer, D. Frincke, and K. Levitt. A template for rapid prototyping of operating systems. In Winkler [Win91], pages 119–127. ISBN 0-8186-2175-3. LCCN QA76.9.C65 I577 1990. IEEE Cat. No.91TH0380-6.

Ahmad:2008:MCC

[Ahm08] Feisal S. Ahmad. MINIX 3 C compiler performance: Comparing the Amsterdam Compiler Kit to the GNU Compiler Collection on x86 systems. Bachelor’s thesis, Department of Computer Science, Faculty of Sciences, Vrije Universiteit, Amsterdam, The Netherlands, June 9, 2008. 27 + 26 (appendix) pp. URL http:

//www.minix3.org/theses/
ahmad-cc-performance.pdf;
 http://www.minix3.org/
theses/ahmad-cc-performance_
appendix.pdf.

Alting:2006:PMP

[Alt06a] Ingmar Alting. MinixPPC: A port of the MINIX OS to the PowerPC platform: Creating a programming model for architecture independency. Master’s thesis in computer science, Department of Computer Science, Faculty of Sciences, Vrije Universiteit, Amsterdam, The Netherlands, September 15, 2006. xii + 83 + 59 (Appendices)

pp.  URL http://www.
minix3.org/theses/alting-
powerpc-port.pdf.

Alting:2006:MPM

[Alt06b] Ingmar A. Alting. MinixPPC: a port of the MINIX OS to the PowerPC platform: Creating a programming model for architecture independency [sic]. Masters thesis, Department of Computer Science, Vrije Universiteit, Amsterdam, The Netherlands, September 15, 2006. xi + 83 + 59 pp.

URL http://www.minix3.
org/doc/alting_thesis.pdf.

Anglin:1991:AMP

[Ang91] Elizabeth Anglin. Addition of a message passing measure to MINIX (a process-based operating system). Thesis (m.s.), Kansas State University, Manhattan, KS, USA, 1991. iii + 90 pp.

Anonymous:1987:ESC

[Ano87] Anonymous, editor. EUUG Spring ’87 Conference Proceedings. EurOpen, Buntingford, Herts, UK, 1987.

Anonymous:1989:SPA

[Ano89a] Anonymous, editor. SCCS Proceedings. 22nd Annual Small College Computing Symposium. Univ. Wisconsin-Eau Claire, Eau Claire, WI, USA, 1989.

Anonymous:1989:UEC

[Ano89b] Anonymous, editor. UNIX:

European challenges: proceedings of the Spring 1989 EUUG conference, April 3– 7, 1989, Palais des Congres, Brussels, Belgium. European UNIX Users Group, Buntingford, Herts, UK, 1989. ISBN 0-9513181-2-8. LCCN QA76.76.O63U54514 1989.

Anonymous:1990:NCS

[Ano90a]  Anonymous, editor.  13th

National Computer Security Conference. Proceedings Information Systems Security. Standards — the Key to the Future. National Institute for Standards and Technology, Gaithersburg, MD, USA, 1990. 2 vol.

Anonymous:1990:AOS

[Ano90b] Anonymous. Alternative operating systems, part 5: Unix with a microscope: Minix, a low-cost Unix, runs on ordinary personal computers. Byte Magazine, 15(13):345– 346, December 1990. CODEN BYTEDJ. ISSN 0360-5280.

Anonymous:1990:ISS

[Ano90c] Anonymous, editor. Information systems security: standards — the key to the future: 13th National Computer Security Conference: Omni Shoreham Hotel, Washington, DC, 1–4 October, 1990: proceedings. National Institute of Standards and Technology, National Computer Security Center, Gaithersburg, MD, USA, 1990.

Anonymous:1990:MCL

[Ano90d]  Anonymous.  MINIX: a

‘CLONIX’ for learning how to build operating systems. Novatica, 16(86):79–82, 1990.

CODEN NOVAEC. ISSN 0211-2124.

Anonymous:1992:EUG

[Ano92] Anonymous, editor. EurOpen.

UNIX Goes East. Proceedings of the Autumn 1991 EurOpen Conference. EurOpen, Buntingford, Herts, UK, 1992.

Anonymous:1995:SCC

[Ano95] Anonymous, editor. Small College computing: Annual symposium; 28th — April 1995, Sioux Falls, SD, SCCS -PROCEEDINGS- 1995; 28th. SCCS, ????, 1995.

Anonymous:2017:MWU

[Ano17] Anonymous. The most widely used OS in the world is the least known. Web site, November 6, 2017. URL https://

liveatpc.com/widely-used-
os-world-least-known/.

Appuswamy:2014:BFB

[App14]  Raja Appuswamy.  Build-

ing a File-Based Storage Stack: Modularity and Flexibility in Loris. Ph.D. thesis, ????, ????, ???? 2014.

????  pp.  URL http:

//www.minx3.org/theses/
Raja_Appuswamy_PhD_thesis.
pdf.

Ashton:1997:UIN

[Ash97] Paul Ashton. Using interaction networks for visualization of message passing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):272– 276, March 1997. CODEN SIGSD3. ISBN 0-89791-889-4. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
1997.bib;   ftp://ftp.math.
utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/
Os/minix.bib.

Appuswamy:2013:FLH

[AvMST13] Raja Appuswamy, David C. van Moolenbroek, Sharan Santhanam, and Andrew S. Tanenbaum. File-level, hostside flash caching with Loris. In IEEE, editor, Proceedings of the 19th International Conference on Parallel and Distributed Systems (ICPADS ’13), Seoul, Korea, December 2013, pages 34–42. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN 1-4799-2082-7. LCCN ???? URL http://

www.minix3.org/docs/conf/
icpads-2013.pdf.

Appuswamy:2010:BLR

[AvMT10a] Raja Appuswamy, David C. van Moolenbroek, and Andrew S. Tanenbaum. Blocklevel RAID is dead. In ????, editor, Proceedings of the Second Workshop on Hot Topics in Storage and File Systems (HotStorage ’10), Boston, June 2010, page ?? ACM Press, New York, NY 10036, USA, 2010. ISBN ???? LCCN ????

URL http://www.minix3.
org/docs/loris/HotStorage2010.
pdf.

Appuswamy:2010:LDM

[AvMT10b] Raja Appuswamy, David C.

van Moolenbroek, and Andrew S. Tanenbaum. Loris

— a dependable, modular file-based storage stack. In IEEE, editor, Proceedings 16th IEEE Pacific Rim International Symposium on Dependable Computing: PRDC 2010: 13–15 December 2010, Tokyo, Japan, pages 165–

174. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-8975-X, 0-7695-4289-

1. LCCN QA76.9.F38. URL

http://ieeexplore.ieee.
org/servlet/opac?punumber=
5702592;   http://www.minix3.
org/docs/loris/PRDC2010.
pdf.

Appuswamy:2011:FMF

[AvMT11] Raja Appuswamy, David C.

van Moolenbroek, and Andrew S. Tanenbaum. Flexible, modular file volume virtualization in Loris. In André Brinkmann and David Pease, editors, Proceedings of the 27th IEEE Conference on Mass Storage Systems and Technologies (MSST ’11), 23–27 May 2011, Denver, CO, USA, pages 1– 14. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4577-0428-5, 1-4577-0427-7. LCCN TK7895.M4. URL

                                        http://ieeexplore.ieee.
                                        org/servlet/opac?punumber=
                                        5910685;   http://www.minix3.
                                        org/docs/loris/MSST2011.
                                        pdf.

Appuswamy:2012:IFB

[AvMT12] Raja Appuswamy, David C. van Moolenbroek, and Andrew S. Tanenbaum. Integrating flash-based SSDs into the storage stack. In ????, editor, Proceedings of the 28th IEEE Conference on Mass Storage Systems and Technologies (MSST ’12), 16–20 April 2012, Pacific Grove, CA, pages 1–

12. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-1747-0. LCCN ????

                                        URL http://www.minix3.
                                        org/docs/loris/MSST2012.
                                        pdf.

Appuswamy:2014:CCE

[AvMT14] Raja Appuswamy, David C. van Moolenbroek, and Andrew S. Tanenbaum. Cache, cache everywhere, flushing all hits down the sink: On exclusivity in multilevel, hybrid caches. In ????, editor, 29th IEEE Symposium on Massive Storage Systems and Technologies: (MSST2013): Research Track: in Long Beach, California: [May 6, 2013–May 10, 2013], pages 1–14. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2014. ISBN 1-4799-0218-7. LCCN TK7895.M4. URL

http://ieeexplore.ieee.
org/servlet/opac?punumber=
6554307;   http://www.minix3.
org/docs/conf/msst-2013.
pdf.

btarunr:2017:ICC

[bta17]  btarunr.  Intel CPU on-

chip management engine runs on MINIX. Web site, November 5, 2017. URL

https://www.techpowerup.
com/238514/intel-cpu-on-
chip-management-engine-
runs-on-minix.

Byfield:2010:IM

[Byf10] Bruce Byfield. An introduction to MINIX. Linux Journal, 2010(194):5:1–5:??, June 2010. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Chappelow:1990:DIS

[CAH90] Stephen W. Chappelow, Steven F.

Ackerman, and Stephen J. Hartley. Design and implementation of a swapper for the MINIX operating system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22 (4):55–59, December 1990.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ciancarini:1993:LMM

[CG93] P. Ciancarini and N. Guerrini. Linda meets Minix. Operating Systems Review, 27(4): 76–92, October 1993. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chittoor:1995:ISS

[Chi95] Chandra S. Chittoor. Implementation of sockets on SunOS Minix. Thesis (m.s.), Kansas State University, Manhattan, KS, USA, 1995. iii + 75 pp.

Christie:1996:DAK

[Chr96] Dave Christie. Developing the AMD-K5 architecture: Flying without instruments: the independent development on the x86 processor. IEEE Micro, 16 (2):16–26, March/April 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VII, Stanford University, Stanford, California, August 1995.

Cox:2001:PX

[Cox01] David Cox. Parsing XML. Dr. Dobb’s Journal of Software Tools, 26(1):96, 98, 100, January 2001. CODEN DDJOEB. ISSN 1044-789X. URL

                                        http://www.ddj.com/ftp/
                                        2001/2001_01/minixml.txt;

http://www.ddj.com/ftp/
2001/2001_01/minixml.zip.

Cushing:1988:IMC

[Cus88] David Bruce Cushing. The implementation of multicast communication in the MINIX operating system kernel. Thesis (M.Comp.Sc.), School of Computer Science, Technical University of Nova Scotia, Halifax, NS, Canada, 1988. viii + 119 pp.

deJonge:1993:LDN

[dJKH93] Wiebren de Jonge, M. Frans Kaashoek, and Wilson C. Hsieh. The Logical Disk: a new approach to improving file systems. Operating Systems Review, 27(5):15–28, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). 14th ACM Symposium on Operating Systems Principles, Ashville, NC, USA.

deSmit:2008:FHS

[dS08] Jens de Smit. Finding hot spots in MINIX 3. Bachelor’s thesis, Department of Computer Science, Faculty of Sciences, Vrije Universiteit, Amsterdam, The Netherlands, January 7, 2008. 15 pp.

URL http://www.minix3.
org/theses/smit-finding-
hot-spots.pdf.

Du:2006:NAC

[DSX06] Wenliang Du, Mingdong Shang, and Haizhi Xu. A novel approach for computer security education using Minix instructional operating system. Computers & Security, 25(3):190–200, May 2006. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167404805001550.

Donaldson:1990:TMW

[DTC90] A. L. Donaldson, J. W. Taylor, Jr., and D. M. Chizmadia. Trusted MINIX: a worked example. In Anonymous [Ano90a], pages 307–317 (vol. 1). 2 vol.

Du:2007:SSI

[DTW07] Wenliang Du, Zhouxuan Teng, and Ronghua Wang. SEED: a suite of instructional laboratories for computer SEcurity EDucation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):486–490, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Durr:1989:MAS

[Dur89] C. L. Durr. Multis for the Atari ST (multi-user packages). Chip, July 1989. CO-DEN CHIPDP. ISSN 0170-6632.

Du:2008:SSI

[DW08] Wenliang Du and Ronghua Wang. SEED: a suite of instructional laboratories for computer security education. ACM Journal on Educational Resources in Computing (JERIC), 8(1):3:1–3:??, March 2008. CODEN ???? ISSN 1531-4278.

Ermolov:2017:DIM

[EG17] Mark Ermolov and Maxim Goryachy. Disabling Intel ME 11 via undocumented mode. Web site, August 28,

2017. URL http://blog.
ptsecurity.com/2017/08/
disabling-intel-me.html.

Fugini:1988:EUP

[FBM88] M. G. Fugini, R. Bellinzona, and G. Martella. An extension to Unix protection mechanisms to support flexible resource sharing and discretionary authorization. In IEEE [IEE88a], pages 663– 671.

Ference:1991:IMC

[Fer91] James M. Ference. Implementation of Minix on the CT-MiniFrame. Thesis (m.s.), San Francisco State University, San Francisco, CA, USA, 1991. xi + 380 pp.

Fraser:2006:AFS

[FPA06] Timothy Fraser, Nick L.

Petroni, Jr., and William A. Arbaugh. Applying flowsensitive CQUAL to verify MINIX authorization check placement: 3. In ACM [ACM06b], pages 3–6. ISBN 1-59593-374-3. LCCN ????

Frankl:2002:IPA

[Fra02] Phyllis G. Frankl, editor. IS-STA 2002: proceedings of the ACM SIGSOFT International Symposium on Software Testing and Analysis, Roma, Italy, July 22–24, 2002, volume 27(4) of Software engineering notes. ACM Press, New York, NY 10036, USA, 2002. ISBN 1-58113-562-9. LCCN ????

Frazelle:2019:OSF

[Fra19] Jesse Frazelle. Open source firmware. Communications of the ACM, 62(10):34–38, October 2019. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:

                                        //cacm.acm.org/magazines/
                                        2019/10/239673/fulltext.

Fresquez:1990:SIM

[Fre90] Vicente Fresquez. A swapping implementation for the MINIX operating system. Thesis (m.s.), University of Texas at El Paso, El Paso, TX, USA, 1990. viii + 122 pp.

Giuffrida:2010:WCN

[GCT10] Cristiano Giuffrida, Lorenzo Cavallaro, and Andrew S. Tanenbaum. We crashed, now what? In ????, editor, Proceedings of the Sixth Workshop on Hot Topics in System Dependability (HotDep ’10), Vancouver, BC, Canada, October 2010, pages 1–8. USENIX Association, Berkeley, CA, USA, 2010. ISBN ???? LCCN QA76.9.A25.

URL http://www.cs.vu.nl/
%7Egiuffrida/papers/hotdep-
2010.pdf.

Giuffrida:2013:PAV

[GCT13] Cristiano Giuffrida, Lorenzo Cavallaro, and Andrew S. Tanenbaum. Practical automated vulnerability monitoring using program state invariants. In IEEE, editor, Proceedings of the 43rd International Conference on Dependable Systems and Networks (DSN ’13), Budapest, Hungary, 24– 27 June 2013, pages 1–12. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN 1-4673-6471-1. ISSN 1530-0889. LCCN QA76.9.F38 I58 2013. URL

http://ieeexplore.ieee.
org/xpl/mostRecentIssue.
jsp?punumber=6569391;
http://www.minix3.org/docs/
conf/dsn-2013.pdf.

Gull:1989:MMH

[GD89a] A. Gull and S. K. Das.

Memory management hardware: panacea or pain? In Anonymous [Ano89b], pages 217–221. ISBN 0-9513181-2-8. LCCN QA76.76.O63U54514 1989.

Gull:1989:PMO

[GD89b] Aarron Gull and Sunil K. Das.

A port of the MINIX operating system to the Atari ST. European UNIX Systems User Group, EUUG Newsletter, 9 (1):2–14, Spring 1989. CO-DEN EONLE8. ISSN 1011-4211.

Gerofi:2006:DIM

[Ger06a] Balázs Gerőfi. Design and implementation of the MINIX virtual file system. Master’s thesis in computer science, Department of Computer Science, Faculty of Sciences, Vrije Universiteit, Amsterdam, The Netherlands, August 2006. xii

                                        + 108 pp. URL http://www.
                                        minix3.org/theses/gerofi-
                                        minix-vfs.pdf.

Gerofi:2006:MVD

[Ger06b] Balázs Gerőfi. MINIX VFS: Design and implementation of the MINIX Virtual File system. Masters thesis, Department of Computer Science, Vrije Universiteit, Amsterdam, The Netherlands, August 2006. x + 108 pp.

                                        URL http://www.minix3.
                                        org/doc/gerofi_thesis.pdf.

Gammill:1989:DFS

[GH89] R. Gammill and J. Hernes. A distributed file system for MINIX. In Anonymous [Ano89a], pages 151–160.

Giuffrida:2013:BFF

[GIKT13] Cristiano Giuffrida, Călin Iorgulescu, Anton Kuijsten, and Andrew S. Tanenbaum. Back to the future: Faulttolerant live update with timetraveling state transfer. In ????, editor, Proceedings of the Large Installation System Administration Conference (LISA ’13), Washington, DC, November 2013, pages 89–104. USENIX Association, Berkeley, CA, USA, 2013. ISBN ???? LCCN ????

URL http://www.minix3.
org/docs/conf/lisa-2013.

pdf. Best student paper award.

Giuffrida:2014:SAL

[Giu14]  Cristiano Giuffrida.  Safe

and Automatic Live Update. Ph.D. thesis, ????, ????, ???? 2014. ???? pp. URL http:

//www.minix3.org/theses/
Cristiano_Giuffrida_PhD_
thesis.pdf.

Giuffrida:2012:EOS

[GKT12] Cristiano Giuffrida, Anton Kuijsten, and Andrew S. Tanenbaum. Enhanced operating system security through efficient and fine-grained address space randomization. In ????, editor, Proceedings of the 21st USENIX Security Symposium (USENIX Sec ’12), Bellevue, WA, Aug, 2012, pages 40–55. USENIX Association, Berkeley, CA, USA, 2012. ISBN ???? LCCN ????

URL http://www.cs.vu.nl/
%7Egiuffrida/papers/usenixsec-
2012.pdf.

Giuffrida:2013:EDF

[GKT13a] Cristiano Giuffrida, Anton Kuijsten, and Andrew S.

Tanenbaum. EDFI: A dependable fault injection tool of dependability benchmarking experiments. In ????, editor, Proceedings of the Pacific Rim International Symposium of Dependable Computing (PRDC ’13), Vancouver, BC, Canada, December 2013, pages 31–40. ????, ????, 2013. ISBN ???? LCCN ???? URL http://

www.minix3.org/docs/conf/
prdc-2013.pdf.

Giuffrida:2013:SAL

[GKT13b] Cristiano Giuffrida, Anton Kuijsten, and Andrew S. Tanenbaum. Safe and automatic live update for operating systems. In Vivek Sarkar and Rastislav Bodik, editors, AS-PLOS XVIII: eighteenth International Conference on Architectural Support for Programming Languages and Operating Systems: March 16– 20, 2013, Houston, Texas, USA, pages 279–292. ACM Press, New York, NY 10036, USA, 2013. ISBN 1-4503-1870-3. LCCN QA76.9.A73.

URL http://dl.acm.org/
citation.cfm?id=2451116;
 http://www.minix3.org/
docs/conf/asplos-2013.pdf.

Guerrero:1993:IEA

[GLG93] R. Guerrero, L. Leguizamon, and R. Gallard. Implementation and evaluation of alternative process schedulers in MINIX. Operating Systems Review, 27(1):79–100, January 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Grehan:1990:MMA

[Gre90] Rick Grehan. Multitasking for the masses: An analysis of different tools to put multitasking on your desk with just a PC or a Macintosh. Byte Magazine, 15(2):279–280, 282, 284, 286, 288, 334, February 1990. CODEN BYTEDJ. ISSN 0360-5280.

Giuffrida:2009:CUN

[GT09] Cristiano Giuffrida and Andrew S. Tanenbaum. Cooperative update: a new model for dependable live update. In Shail Arora, editor, HOTSWUP ’09: Proceedings of the second International Workshop on Hot Topics in Software Upgrades, October 25, 2009, Orlando, Florida, pages 1–6. ACM Press, New York, NY 10036, USA, 2009. ISBN 1-60558-723-0. LCCN QA76.76.S64 I584 2009. URL http://dl.

acm.org/citation.cfm?id=
1656437;   http://www.cs.
vu.nl/%7Egiuffrida/papers/
hotswup-2009.pdf.

Giuffrida:2012:SAS

[GT12] Cristiano Giuffrida and Andrew S. Tanenbaum. Safe and automated state transfer for secure and reliable update. In IEEE, editor, Proceedings of the Fourth International Workshop on Hot Topics in Software Upgrades (HotSWUp ’12), Zürich, 3 June 2012. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-1764-0, 1-4673-1765-9. LCCN QA76.76.S64.

                                        URL http://www.minix3.
                                        org/docs/workshop/hotswup-
                                        2012.pdf.

Guha:1989:EMO

[Guh89] Amitava Guha. Enhancement of Minix operating system. Thesis (m.s.), Department of Computer Science, Southern Illinois University at Carbondale, Carbondale, IL, USA, 1989. 44 pp.

Hartley:1990:EMO

[Har90] Stephen J. Hartley. Experience with MINIX in an operating systems lab. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(3):34– 38, September 1990. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hays:1989:OSC

[Hay89] James H. Hays. An operating systems course using Minix. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 21 (4):11–12, December 1, 1989. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Herder:2006:WAD

[HBG^{+}06a] J. N. Herder, H. Bos, B. Gras, P. Homburg, and A. S. Tanenbaum. Who’s afraid of dead device drivers? Technical Report IR-CS-D29, Department of Computer Science, Vrije Universiteit, Amsterdam, The Netherlands, 2006. ?? pp.

Herder:2006:AFR

[HBG^{+}06b] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. The architecture of a faultresilient operating system. In ????, editor, Proceedings of the 12th ASCI Conference (ASCI ’06), Lommel, Belgium, June 2006, pages 74–81. ????, ????, 2006. ISBN ???? LCCN

????  URL http://www.
minix3.org/docs/jorrit-
herder/asci06.pdf.

Herder:2006:CHD

[HBG^{+}06c] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. Construction of a highly dependable operating system. In IEEE, editor, Proceedings, 6th European Dependable Computing Conference EDCC 2006: 18–20 October 2006, Coimbra, Portugal, pages 3–12. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 0-7695-2648-9. LCCN QA76.9.F38. URL

http://ieeexplore.ieee.
org/servlet/opac?punumber=
4020815;   http://www.eu-
egee.org/egee_events/events/
edcc-6-sixth-european-dependable-
computing-conference-18-
20-october-2006-coimbra-
portugal/;   http://www.
minix3.org/doc/EDCC-2006.
pdf;   http://www.minix3.
org/docs/jorrit-herder/
edcc06.pdf.

Herder:2006:MHR

[HBG^{+}06d] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. MINIX 3: a highly reliable, self-repairing operating system. Operating Systems Review, 40(3): 80–89, July 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (elec-

tronic). URL http://www.
minix3.org/doc/OSR-2006.
pdf;   http://www.minix3.
org/docs/jorrit-herder/
osr-jul06.pdf.

Herder:2006:MSP

[HBG^{+}06e] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. Modular system programming in MINIX 3. ;login: the USENIX Association newsletter, 31(2):19– 28, April 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.minix3.

org/docs/login-2006.pdf;
http://www.usenix.org/
publications/login/2006-
04/openpdfs/herder.pdf.

Herder:2006:RUR

[HBG^{+}06f] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. Reorganizing UNIX for reliability. In Jesshope and Egan [JE06], pages 81– 94. ISBN 3-540-40056-7, 3-540-40058-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A73 A28 2006. URL http://

www.minix3.org/doc/ACSAC-
2006.pdf;   http://www.
minix3.org/docs/jorrit-
herder/acsac06.pdf.

Herder:2007:FRD

[HBG^{+}07a] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. Failure resilience for device drivers. In IEEE, editor, Proceedings of the 37th Annual IEEE/IFIP International Conference on Dependable Systems and Networks (DSN ’07), Edinburgh, UK, June 2007, pages 41–50. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2007. ISBN 0-7695-2855-4. LCCN QA76.9.F38 I324

2007.  URL http://www.
minix3.org/docs/jorrit-
herder/dsn-dccs07.pdf.

Herder:2007:RFR

[HBG^{+}07b] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. Roadmap to a failureresilient operating system. ;login: the USENIX Association newsletter, 32(1):14– 20, February 2007. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.minix3.
org/docs/login-2007.pdf;
 https://www.usenix.org/
publications/login/february-
2007-volume-32-number-1/
roadmap-failure-resilient-
operating-system.

Herder:2008:CIT

[HBG^{+}08] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. Countering IPC threats in multiserver operating systems. In IEEE, editor, PRDC 2008: the 14th IEEE Pacific Rim International Symposium on Dependable Computing: proceedings: 15–17 December, 2008, Taipei, Taiwan, pages 112–121. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2008. ISBN 0-7695-3448-1, 1-4244-4094-7. LCCN QA76.9.F38 P33 2008. URL

http://ieeexplore.ieee.
org/servlet/opac?punumber=
4725263;   http://www.minix3.
org/docs/jorrit-herder/
prdc08.pdf.

Herder:2009:FID

[HBG^{+}09] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. Fault isolation for device drivers. In IEEE, editor, Proceedings of the 39th IEEE/ IFIP International Conference on Dependable Systems and Networks (DSN 09), June 29, 2009–July 2, 2009, Estoril, Lisbon, Portugal, pages 33–42. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 1-4244-4421-7, 1-4244-4422-5 (print). LCCN QA76.9.F38.

URL http://dl.acm.org/
citation.cfm?id=2451116;
 http://ieeexplore.ieee.
org/servlet/opac?punumber=
5243642;   http://www.minix3.
org/docs/jorrit-herder/
dsn-dccs09.pdf. IEEE cat-

alog number CFP09048-CDR.

Herder:2006:LMB

[HBT06] Jorrit N. Herder, Herbert Bos, and Andrew S. Tanenbaum. A lightweight method for building reliable operating systems despite unreliable device drivers. Technical Report IR-CS-018, Department of Computer Science, Vrije Universiteit, Amsterdam, The Netherlands, January 2006.

14 pp. URL http://www.
minix3.org/doc/reliable-
os.pdf.

Hruby:2013:HMW

[HBT13a] Tomáš Hrubý, Herbert Bos, and Andrew S. Tanenbaum. Heterogeneous multicores: When slower is faster. In ????, editor, Proceedings of the Third Workshop on Systems for Future Multicore Architectures (SFMA ’13), Prague, April 2013, page ?? ACM Press, New York, NY 10036, USA, 2013. ISBN ???? LCCN

                                        ????  URL http://www.
                                        minix3.org/docs/workshop/
                                        sfma-2013.pdf.

Hruby:2013:WSF

[HBT13b] Tomáš Hrubý, Herbert Bos, and Andrew S. Tanenbaum. When slower is faster: on heterogeneous multicores for reliable systems. In ????, editor, Proceedings of the Annual Technical Conference (ATC ’13), San Jose, CA, June 2013, pages 255–266. USENIX Association, Berkeley, CA, USA, 2013. ISBN ???? LCCN

                                        ????  URL http://www.
                                        minix3.org/docs/conf/atc-
                                        2013.pdf.

Hruby:2014:SMS

[HBT14] Tomáš Hrubý, Herbert Bos, and Andrew S. Tanenbaum. Towards optimal scheduling of multiserver system components. In ????, editor, Proceedings of the Fourth Workshop on Systems for Future Multicore Architectures (SFMA ’14), Amsterdam, April 2014, page ?? ????, ????, 2014. ISBN ???? LCCN

????  URL http://www.
minix3.org/docs/workshop/
sfma-2014.pdf.

Hoffman:1990:FSM

[Hd90] P. K. Hoffman and G. de V.

Smit. A file server for a multi-transputer Unix system. South African Computer Journal, October 1990. ISSN 1015-7999.

Heiser:2005:SES

[Hei05] Gernot Heiser. Secure embedded systems need microkernels. ;login: the USENIX Association newsletter, 30(6): ??, December 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-12/pdfs/heiser.pdf.

Hernes:1989:DFS

[Her89] Jeffrey Lee Hernes. A distributed file server for MINIX. Thesis (m.s.), North Dakota State University, Fargo, ND, USA, 1989. iv + 43 pp.

Her:1990:ISR

[Her90] Sheau-Chuen Her. Implement a simple remote procedure call model in minix. Thesis (m.s.), California State University, Chico, Chico, CA, USA, 1990. viii + 43 pp.

Herder:2005:ROS

[Her05a] Jorrit N. Herder. Reliable operating systems: Research summary. In ????, editor, Proceedings of the First EuroSys Doctoral Workshop, 23 October 2005, Brighton, UK, page ?? ????, ????, 2005. ISBN ???? LCCN

????  URL http://www.
minix3.org/docs/jorrit-
herder/eurosys-dw05-talk.
pdf.

Herder:2005:TTM

[Her05b] Jorrit N. Herder. Towards a true microkernel operating system: A revision of MINIX that brings quality enhancements and strongly reduces the kernel in size by moving device drivers to userspace. Master of Science thesis, Department of Computer Science, Faculty of Sciences, Vrije Universiteit, Amsterdam, The Netherlands, February 23, 2005. xvi + 117 pp.

URL http://www.minix3.
org/doc/herder_thesis.pdf;
 http://www.minix3.org/
theses/herder-true-microkernel.
pdf.

Herder:2006:RDD

[Her06] Jorrit N. Herder. Reincarnation of dead device drivers. In ????, editor, Proceedings of the First EuroSys Authoring Workshop, Leuven, Belgium, April 2006, page ?? ????, ????, 2006. ISBN ???? LCCN

????  URL http://www.
minix3.org/docs/jorrit-
herder/eurosys-aw06-talk.
pdf.

Herder:2010:BDO

[Her10] Jorrit Niek Herder. Building a Dependable Operating System: Fault Tolerance in MINIX 3. Ph.D. thesis, Vrije Universiteit Amsterdam, Amsterdam, The Netherlands, September 2010. 2 + xvii + 165 pp. URL http:

//www.minix3.org/theses/
herder_phd_thesis.pdf.

Hoffmann:2010:LBQb

[Hof10] Leah Hoffmann. Last byte:

Q&A: Systematic thinking: Andrew S. Tanenbaum talks about MINIX, microkernels, and electronic voting systems. Communications of the ACM, 53(4):112–ff, April 2010. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Howatt:2002:OSP

[How02] James Howatt. Operating systems projects: Minix revisited. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):109–111, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2002.bib.

Hruby:2012:KNW

[HVBT12] Tomáš Hrubý, Dirk Vogt, Herbert Bos, and Andrew S. Tanenbaum. Keep net working — on a dependable and fast networking stack. In IEEE, editor, Proceedings of the 42nd Annual IEEE/IFIP International Conference on Dependable Systems and Networks (DSN ’12), Boston, MA, June 2012, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-1623-7, 1-4673-1624-5. LCCN QA76.9.F38 I58 2012. URL

http://ieeexplore.ieee.
org/servlet/opac?punumber=
6253623;   http://www.minix3.
org/theses/dsn2012.pdf.

Herder:2009:DDF

[HvMA^{+}09] Jorrit N. Herder, David C. van Moolenbroek, Raja Appuswamy, Bingzheng Wu, Ben Gras, and Andrew S. Tanenbaum. Dealing with driver failures in the storage stack. In IEEE, editor, LADC 2009: 2009 Fourth Latin-American Symposium on Dependable Computing: proceedings: 1–4 September 2009, Joõ Pessoa, Paraíba, Brazil, pages 119–

126. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 0-7695-3760-X, 1-4244-4678-3 (print). LCCN QA76.9.F3.

URL http://www.minix3.
org/docs/jorrit-herder/
ladc09.pdf.  Best paper

award.

IEEE:1988:AIT

[IEE88a] IEEE, editor. Artificial intelligence: theory and applications: proceedings, International Computer Science Conference ’88: 19–21 December, 1988, the Excelsior Hotel, Causeway Bay, Hong Kong. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988.

IEEE:1988:FAC

[IEE88b] IEEE, editor. Fourth Aerospace Computer Security Applications Conference: Orlando, Florida, December 12–16, 1988. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0895-1. LCCN QA76.9.A25 A39 1988. IEEE Cat. No.CH2619-5.

IEEE:1994:SAI

[IEE94] IEEE, editor. System architecture and integration: proceedings of the 20th EUROMICRO Conference, EUROMICRO 94, September 5–8, 1994, Liverpool, England. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6430-4. LCCN QA76.9.A73 E94 1994.

Iorgulescu:2013:SAL

[Ior13] Călin Iorgulescu. Safe and automatic live update for Unix applications. Master’s thesis, ????, ????, ???? 2013.

???? pp. URL http://www.
minix3.org/theses/Calin_

                                        Iorgulescu_Master_Thesis.
                                        pdf.

Jesshope:2006:ACS

[JE06] Chris Jesshope and Colin Egan, editors. Advances in computer systems architecture: 11th Asia-Pacific conference, ACSAC 2006, Shanghai, China, September 6– 8, 2006, proceedings, volume 4186 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-40056-7 (softcover). LCCN QA76.9.A73 A28 2006.

Kachel:1989:MMB

[Kac89] Timothy Paul Kachel. MIN-NET, a MINIX based broadcast network using RS232. Thesis (m.s.), North Dakota State University, Fargo, ND, USA, 1989. iv + 67 pp.

Kanapoulos:1992:SIW

[Kan92] N. Kanapoulos, editor. The Second International Workshop on Rapid System Prototyping: Research Triangle Park, North Carolina, USA, June 11–13, 1991: shortening the path from specification to prototype. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-3040-X. LCCN QA76.76.D47 I598 1991. IEEE Catalog No. 92TH0454-9.

Kelly:2006:PMX

[Kel06] Ivan Kelly. Porting MINIX to Xen. Final year project, Department of Computer Science, University of Limerick, Limerick, Ireland, May 8, 2006. vii + 45 pp. URL http://

minixonxen.skynet.ie/cgi-
bin/trac.cgi/attachment/
wiki/Report/Report.pdf?
format=raw; http://www.
minix3.org/theses/kelly-
xen.pdf.

Koo:1988:SSM

[KK88] Yong Wan Koo and Young Chan Kim. A study on the scheduling mechanism for real time system. Journal of the Korea Information Science Society = Chongbo Kwahakhoe nonmunji, 15(3):158–170, June 1988. CODEN HJKHDC. ISSN 0258-9125.

Kobylanski:1989:IPS

[Kob89] Stanley George Kobylanski.

An implementation of process swapping in MINIX (a message passing oriented operating system). Thesis (m.s.), Kansas State University, Manhattan, KS, USA, 1989. vi + 63 + 59 pp.

Koch:1990:MTS

[Koc90] L. Koch. MINIX/THL a training system for instruction in operating systems. Informatik, Informationen Reporte, 1990. ISSN 0233-2582.

Kavka:1993:EDM

[KPG93] C. Kavka, M. Printista, and R. Gallard. Extending device management in Minix. Operating Systems Review, 27 (2):35–43, April 1993. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kuijsten:2012:POS

[Kui12] Anton Kuijsten. Polymorphic operating systems. Master’s thesis, ????, ????, ???? 2012. ???? pp. URL http:

                                        //www.minix3.org/theses/
                                        kuijsten-polymorphic-os.
                                        pdf.

Lakshmi:2004:IMM

[Lak04] Narasimhan Krishnan Lakshmi. Improving MINIX memory management. Thesis (M.S.), Kansas State University, Manhattan, KS 66506, USA, 2004. vi + 62 pp.

Larribeau:1990:MDP

[Lar90] Scott Larribeau. The MINIX development project. Thesis (m.s.), California Polytechnic State University, San Luis Obispo, CA, USA, 1990. vi + 145 pp.

Chang:1995:IWS

[lC95] Chiu liang Chang. The implementation of a window system for MINIX 1.3. Thesis (m.s.), California State University, Chico, Chico, CA, USA, 1995. x + 65 pp.

Li:1988:SUI

[LG88] Lianzhi Li and Fushun Guo.

The structure and userinterface of MINIX. Mini-Micro Systems, 9(10):7–10, 15, 1988. CODEN XWJXEH. ISSN 0364-9342.

Li:1993:PMV

[Li93] Xiaohong Li. Porting MINIX to VM. Thesis (m.s.), Teikyo Marycrest University, Tokyo, Japan (??), 1993. v + 83 pp.

Linnenbank:2009:IIP

[Lin09] Niek Q. Linnenbank. Implementing the Intel Pro/1000 on MINIX 3. Course project, Department of Computer Science, Faculty of Sciences, Vrije Universiteit, Amsterdam, The Netherlands, December 12, 2009. i + 41 pp. URL http:

//www.minix3.org/theses/
linnenbank-ipa.pdf.

Linnenbank:2011:IMS

[Lin11] Niek Q. Linnenbank. Implementing Minix on the single chip cloud computer. Master’s thesis, ????, ????, ???? 2011. ???? pp.

URL http://www.minix3.
org/docs/Implementing.MINIX.
On.The.Single.Chip.Cloud.
Computer.pdf.

Louboutin:1992:IPM

[Lou92] S. R. Y. Louboutin. An implementation of a process migration mechanism using Minix. In Anonymous [Ano92], pages 213–224.

Lunduke:2017:PBT

[Lun17a] Bryan Lunduke. Het populairste besturingssysteem ter wereld is … MINIX? (Dutch) [The most popular operating system in the world is … MINIX?]. Web site, November 3, 2017. URL http:

//webwereld.nl/security/
101772-het-populairste-
besturingssysteem-ter-wereld-
is---minix.

Lunduke:2017:MMP

[Lun17b] Bryan Lunduke. MINIX — the most popular OS in the world, thanks to Intel. You might not know it, but inside your Intel system, you have an operating system running in addition to your main OS, MINIX. and it’s raising eyebrows and concerns. Network World, ??(??):??, November 2, 2017. URL https:

//www.networkworld.com/
article/3236064/servers/
minix-the-most-popular-
os-in-the-world-thanks-
to-intel.html.

Maginnis:1988:DCT

[Mag88] P. Tobin Maginnis. Design considerations for the transformation of MINIX into a distributed operating system. In ACM [ACM88], pages 608–

615. ISBN 0-89791-260-8. LCCN QA 76.758 A26 1988.

McGregor:1997:BBD

[McG97] Anthony James McGregor. Block-Based Distributed File Systems. Doctor of Philosophy thesis, University of Waikato, Waikato, New Zealand, ???? 1997. ???? pp.

URL http://hdl.handle.
net/10289/2607;   http:/
/researchcommons.waikato.
ac.nz/bitstream/handle/
10289/2607/thesis.pdf.

Meier:1991:PM

[Mei91] Harald Meier. Portierung von Minix auf 68000-rechner. Master’s thesis, Technische Universität Braunschweig, Braunschweig, Germany, 1991.

Meurs:2006:BPM

[Meu06] Rogier Meurs. Building performance measurement tools for the MINIX 3 operating system. Master’s thesis, Department of Computer Science, Vrije Universiteit, Amsterdam, The Netherlands, August 2006. iv + 59 pp. URL http://

www.minix3.org/doc/meurs_
thesis.pdf;   http://www.
minix3.org/theses/meurs-
profiling-tools.pdf.

Mancina:2009:EDM

[MFH^{+}09] Antonio Mancina, Dario Faggioli, Giuseppe Lipari Jorrit N. Herder, Ben Gras, and Andrew S. Tanenbaum. Enhancing a dependable multiserver operating system with temporal protection via resource reservations. Real-Time Systems, 43(2):177–210, October 2009. CODEN RESYE9.

ISSN 0922-6443 (print), 1573-1383 (electronic). URL

                                        http://link.springer.com/
                                        article/10.1007%2Fs11241-
                                        009-9086-5;   http://www.
                                        minix3.org/docs/jorrit-
                                        herder/rtns08.pdf. Best pa-

per award.

Matsumoto:1995:CSN

[MHY^{+}95] H. Matsumoto, Y. Hashino, H. Yashiro, N. Shinohara, and H. Omura. Computer simulation on nonlinear interaction of intense microwaves with space plasmas. Transactions of the Institute of Electronics, Information and Communication Engineers B-II, J78B-II (3):119–129, March 1995. CO-DEN DTBTEU.

Mull:1991:EST

[MM91] Allison J. Mull and P. Tobin Maginnis. Evolutionary steps toward a distributed operating system: theory and implementation. Operating Systems Review, 25(4):4–13, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Meumann:1990:MDD

[MR90] M. D. Meumann and M. H. Rennhackkamp. MINIX for a distributed database system. South African Computer Journal, October 1990. ISSN 1015-7999.

Naniwadekar:1988:IRM

[Nan88a] Devendra Naniwadekar. Implementation of a resource manager for distributed MINIX. In ACM [ACM88], page 686. ISBN 0-89791-260-8. LCCN QA 76.758 A26 1988.

Naniwadekar:1988:ACD

[Nan88b] Devendra Vithal Naniwadekar.

Approximations to creating a distributed MINIX operating system. Thesis (m.s.), University of Mississippi, Oxford, MS, USA, 1988. v + 89 pp.

Navaux:1988:MA

[NCCN88] P. O. A. Navaux, T. S. Cirano, A. S. Carissimi, and J. C. Netto. M3P-project architecture. In Silveira [Sil88], pages 234–244.

Nakao:1993:ACD

[NKN93] Zensho Nakao, Masaya Kinjo, and Masahiro Nakama. Ada and C: differences as the language for system programming. ACM SIGADA Ada Letters, 13(5):22–31, September/October 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Noll:2004:MDD

[Nol04] Ryan P. Noll. Minix device driver for the AMD PCnet32. Thesis (B.S.), California Polytechnic State University, San Luis Obispo, CA, USA, 2004. various pp.

ONeil:1995:GDD

[OK95] T. E. O’Neil and B. Knudson. A generic device driver for the MINIX operating system. In Anonymous [Ano95], pages 187–193.

Olabe:1997:OSD

[Ola97] M. A. Olabe. Operating systems design and implementation: an integrated approach using Minix. Computers in education journal, 7(1):59–65, January/March 1997. CO-DEN CEJOE7. ISSN 1069-3769.

Ostrand:2002:DFL

[OW02] T. J. Ostrand and E. J. Weyuker. The distribution of faults in a large industrial software system. In Frankl [Fra02], pages 55–64. ISBN 1-58113-562-9. LCCN ????

Pazzini:1992:TMT

[PN92] M. Pazzini and P. Navaux. TRIX, a multiprocessor transputerbased operating system. In Valero et al. [VOJ^{+}92], pages 621–630 (vol. 1). ISBN 84-87867-13-8. LCCN ???? Two volumes.

Priescu:2012:EUH

[Pri12] Valentin Gabriel Priescu. Efficient use of heterogeneous multicore architectures in reliable multiserver systems. Master’s thesis, ????, ????, ???? 2012. ???? pp. URL http:

//www.minix3.org/theses/
priescu_thesis_2012.pdf.

Ramesh:1988:DDM

[Ram88] K. S. Ramesh. Design and development of MINIX distributed operating system. In ACM [ACM88], page 685. ISBN 0-89791-260-8. LCCN QA 76.758 A26 1988.

Rennhackkamp:1990:NDD

[Ren90] M. H. Rennhackkamp. The NRDNIX distributed database management system. South African Computer Journal, January 1990. ISSN 1015-7999.

Roskos:1988:MSP

[Ros88] J. Eric Roskos. MINIX security policy model. In IEEE [IEE88b], pages 393– 399. ISBN 0-8186-0895-1. LCCN ???? Available from IEEE Service Cent (catalog no. 88CH2619-5). Piscataway, NJ, USA.

Rong:1993:LMM

[RT93] Tsai Shang Rong and Lian-Jou Tsai. A logical machine monitor supporting an environment for development and execution of operating systems. The Journal of Systems and Software, 21(1):27– 39, April 1993. CODEN JS-SODM. ISSN 0164-1212.

Swift:2004:RDD

[SABL04] M. M. Swift, M. Annamalai, B. N. Bershad, and H. M. Levy. Recovering device drivers. In ????, editor, Proceedings of the 6th Symposium on Operating System Design and Implementation, pages 1– 15. ACM Press, New York, NY 10036, USA, 2004. ISBN ???? LCCN ????

Sand:1990:IDF

[San90] Erik Martin Sand. Implementation of a distributed file system for MINIX: a paper. Thesis (m.s.), North Dakota State University, ?Fargo, ND, USA, 1990. v + 158 pp.

Severance:2014:ATI

[Sev14] Charles Severance. Andrew S. Tanenbaum: The impact of MINIX. Computer, 47(7):7–8, July 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Silveira:1988:ADV

[Sil88] P. M. Silveira, editor. Anais do VIII Congresso da Sociedade Brasileira de Computacao (VIII Congress of the Brazilian Computing Society). Sociedade Brasileira Computacao, Rio de Janeiro, Brazil, 1988.

Smith:1991:RPU

[Smi91] Richard Smith. Real-time processing under the MINIX system. Thesis (m.sc.), University of Regina, Regina, Saskatchewan, Canada, 1991. 2 microfiches. University Microfilms order no. UMI00319115.

Swift:2010:UMS

[Swi10] Björn Patrick Swift. User mode scheduling in MINIX 3. Individual programming assignment, Department of Computer Science, Faculty of Sciences, Vrije Universiteit, Amsterdam, The Netherlands, October 27, 2010. 20 pp.

URL http://www.minix3.
org/docs/scheduling/report.
pdf.

Szekeres:2011:IDM

[Sze11] Adriana Szekeres. IOMMU driver for MINIX 3. Course project, Department of Computer Science, Faculty of Sciences, Vrije Universiteit, Amsterdam, The Netherlands, May 2011. ii + 18 pp.

URL http://www.minix3.
org/docs/szekeres-iommu.
pdf.

Tanenbaum:2010:MSR

[TAB^{+}10] Andrew S. Tanenbaum, Raja Appuswamy, Herbert Bos, Lorenzo Cavallaro, Cristiano Giuffrida, Jorrit N. Herder, Tomáš Hrubý, Erik van der Kouwe, and David van Moolenbroek. MINIX 3: Status report and current research. ;login: the USENIX Association newsletter, 35(3): 7–13, June 2010. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.minix3.
org/docs/login-2010.pdf;
 https://www.usenix.org/
publications/login/june-
2010-volume-35-number-3/
minix-3-status-report-and-
current-research.

Tanenbaum:1987:OSD

[Tan87a] A. S. Tanenbaum. Operating Systems: Design and Implementation. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1987. ISBN 0-13-637331-3.

Tanenbaum:1987:MAU

[Tan87b] Andrew S. Tanenbaum. MINIX: A UNIX clone with source code for the IBM PC. ;login: the USENIX Association newsletter, 12(2):3–9, March 1987. CODEN LOGNEM. ISSN 1044-6397.

Tanenbaum:1987:MUC

[Tan87c] Andrew S. Tanenbaum. MINIX: a UNIX clone with source code for the IBM PC. ;login: the USENIX Association newsletter, 12(2):3–9, March/April 1987. CODEN LOGNEM. ISSN 1044-6397.

Tanenbaum:1987:MBSa

[Tan87d] Andrew S. Tanenbaum. Minix binaries and sources for 512K IBM PC-AT’s, 1987. 6 computer disks.

Tanenbaum:1987:MBSb

[Tan87e] Andrew S. Tanenbaum. Minix binaries and sources for 640K IBM PC’s, 1987. ISBN 0-13-583873-8. 9 computer disks.

Tanenbaum:1987:UCS

[Tan87f] Andrew S. Tanenbaum. A UNIX clone with source code for operating systems courses. Operating Systems Review, 21 (1):20–29, January 1987. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tanenbaum:1988:MIPb

[Tan88a] Andrew S. Tanenbaum. Minix 1.3 for the IBM PC-AT’s, 1988. ISBN 0-13-583303-5. 5 computer disks.

Tanenbaum:1988:MIPa

[Tan88b] Andrew S. Tanenbaum. MINIX for the IBM PC, XT, and AT. Prentice-Hall software series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1988. ISBN 0-13-584400-2 (paperback). xv + 486 pp. LCCN QA76.76.O63.

Tanenbaum:1988:OSD

[Tan88c] Andrew S. Tanenbaum. Operating System: Design and Implementation. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1988. ISBN 0-13-637331-3.

Tanenbaum:1991:M

[Tan91a] Andrew S. Tanenbaum. MINIX, 1991. 12 computer disks.

Tanenbaum:1991:MIP

[Tan91b] Andrew S. Tanenbaum. MINIX 1.5 5 1/4in for the IBM PC, XT, AT, 386 and PS/2, 1991. ISBN 0-13-585076-2. 17 computer disks.

Tanenbaum:1991:MAS

[Tan91c] Andrew S. Tanenbaum. MINIX 1.5 for the Atari ST and MegaST, 1991. ISBN 0-13-585035-5. 10 computer disks.

Tanenbaum:2001:MOS

[Tan01] Andrew S. Tanenbaum. Modern operating systems. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2001. ISBN 0-13-031358-

0. xxiv + 951 pp. LCCN QA76.76.O63 T359 2001.

Tanenbaum:2015:IM

[Tan15a] Andrew S. Tanenbaum. The impact of MINIX. YouTube interview (10m47s video).,

2015. URL https://www.
youtube.com/watch?app=desktop&
v=86_BkFsb4eIa.

Tanenbaum:2015:M

[Tan15b] Andrew S. Tanenbaum. MINIX 3. Free and Open Source Software Conference (1h3m video), 2015. URL https:/

/www.youtube.com/watch?v=
jiGjp7JHiYs.

Tanenbaum:2015:MRS

[Tan15c] Andrew S. Tanenbaum. MINIX 3: a reliable and secure operating system. Codemotion Rome 2015 conference talk (73m23s video), 2015. URL https:/

/www.youtube.com/watch?v=
jiGjp7JHiYs.

Tanenbaum:2016:LLY

[Tan16] Andrew S. Tanenbaum. Lessons learned from 30 years of MINIX. Communications of the ACM, 59(3):70–78, March 2016. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//cacm.acm.org/magazines/
2016/3/198874/fulltext.

Tanenbaum:2017:OLI

[Tan17] Andrew S. Tanenbaum. An open letter to Intel. Web site, November 2017. URL http://

www.cs.vu.nl/~ast/intel/.

Tsai:1991:ICM

[TC91] Shang Rong Tsai and Ru Jing Chen. Interprocess communication with multicast support in DMINIX operating system. Microprocessing and Microprogramming, 32(1-5):145– 152, August 1991. CODEN MMICDT. ISSN 0165-6074. 17th EUROMICRO Symposium on Microprocessing and Microprogramming. Hardware and Software Design Automation.

Tsai:1994:LBF

[TCJ94] Shang Rong Tsai, Jyh-Tzong Chiou, and Huan-Ting Jen. Load balance facility in distributed MINIX system. In IEEE [IEE94], pages 162–169. ISBN 0-8186-6430-4. LCCN QA76.9.A73 E94 1994.

Tanenbaum:2006:CWM

[THB06] Andrew S. Tanenbaum, Jorrit N. Herder, and Herbert Bos. Can we make operating systems reliable and secure? Computer, 39 (5):44–51, May 2006. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.
computer.org/csdl/mags/

                                        co/2006/05/r5044-abs.html.

Cover feature.

Tiwana:1990:ITI

[Tiw90] Gurumukh Singh Tiwana. Implementation of TCP/IP in the Minix operating system. Thesis (m.s.), Southern Illinois University at Carbondale, Carbondale, IL, USA, 1990. iii + 48 + [1] pp.

Tanenbaum:1992:MRM

[TKS92] Andrew S. (Andrew Stuart) Tanenbaum, Adrie Koolen, and Johan W. Stevenson. Minix 1.5 reference manual. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-579632-6. vii + 709 pp. LCCN QA76.76.O63 M74525 1992.

Tanenbaum:1988:MAS

[TSM88] Andrew S. Tanenbaum, Johan W. Stevenson, and Jost Muller. MINIX for the ATARI ST and MINIX manual for the ATARI ST. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, version 1.1. edition, 1988. ISBN 0-13-584392-8 (disks), 0-13-584434-7 (manual). LCCN QA76.76.O63. 9 computer disks.

Tsai:1993:LMM

[TT93] Shang Rong Tsai and Lian-Jou Tsai. A logical machine monitor supporting an environment for development and execution of operating systems. The Journal of Systems and Software, 21(1):27– 39, April 1993. CODEN JS-SODM. ISSN 0164-1212.

Tanenbaum:1997:OSD

[TW97] Andrew S. Tanenbaum and Albert S. Woodhull. Operating Systems: Design and Implementation. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1997. ISBN 0-13-638677-6. xvii + 939 pp. LCCN QA76.76.O63T36 1997. US$62.00. Includes CD-ROM.

Tanenbaum:2006:OSD

[TW06] Andrew S. Tanenbaum and Albert S. Woodhull. Operating systems: design and implementation. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, third edition, 2006. ISBN 0-13-142938-8. xvii + 1054 pp. LCCN QA76.76.O63 T36 2006.

Tanenbaum:2009:OSD

[TW09] Andrew S. Tanenbaum and Albert S. Woodhull. Operating systems: design and implementation. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, third edition, 2009. ISBN 0-13-505376-5. xvii + 1054 pp. LCCN ????

Tanenbaum:1998:OSS

[TWC98] Andrew S. Tanenbaum, Albert S. Woodhull, and Junko Chigira. Operetingu shisutemu: sekkei to riron oyobi minikkusu ni yoru jisso. Purentisuhoru Shuppan, Tokyo, Japan, second edition, 1998. ISBN 4-89471-047-1. 1033 pp. LCCN ????

Tyson:2017:IME

[Tys17] Mark Tyson. Intel Management Engine runs on MINIX 3 OS. Web site, November 6,

                                        2017. URL https://hexus.
                                        net/tech/news/software/
                                        111857-intel-management-
                                        engine-runs-minix-3-os/.

Vaidyanathan:1996:PMM

[Vai96] Ranjani Vaidyanathan. Porting MINIX to a multiprocessor platform. Thesis (m. s.), Southwest Texas State University, San Marcos, TX, US, 1996. 156 pp.

vanderKouwe:2009:PQV

[vdK09] Erik van der Kouwe. Porting the QEMU virtualization software to MINIX

3. Master’s thesis, ????,

????, ????  2009.  ????

                                        pp. URL http://www.few.
                                        vu.nl/%7Evdkouwe/doc/msc-
                                        thesis-cs-presentation-
                                        erik-van-der-kouwe.pdf;
                                        http://www.minix3.org/theses/
                                        kouwe-qemu.pdf.

vanderKouwe:2014:EDF

[vdKGT14a] Erik van der Kouwe, Cristiano Giuffrida, and Andrew S. Tanenbaum. Evaluating distortion in fault-injection experiments. In IEEE, editor, HASE 2014: 2014 IEEE 15th International Symposium on High-Assurance Systems Engineering: proceedings, 9– 11 January, 2014, Miami, Florida, USA, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2014. ISBN 1-4799-3465-8 (print). LCCN TA168. URL

http://ieeexplore.ieee.
org/servlet/opac?punumber=
6754245;   http://www.minix3.
org/docs/conf/hase-2014.

pdf. Best paper award.

vanderKouwe:2014:SSI

[vdKGT14b] Erik van der Kouwe, Cristiano Giuffrida, and Andrew S. Tanenbaum. On the soundness of silence: Investigating silent failures using fault injection experiments. In IEEE, editor, Proceedings 2014 Tenth European Dependable Computing Conference: EDCC 2014: Newcastle upon Tyne, Newcastle, United Kingdom: 13– 16 May 2014, pages 118–129. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2014. ISBN 1-4799-3803-3, 1-4799-3804-1. LCCN QA76.9.F38. URL http://

www.minix3.org/docs/conf/
edcc-2014.pdf.

Veerman:2009:DUF

[Vee09] Thomas Veerman. Dynamic updates and failure resistance for the Minix file server. Master of Science’s thesis, Department of Computer Science, Faculty of Sciences, Vrije Universiteit, Amsterdam, The Netherlands, May 11, 2009. ix + 71 pp. URL http:

                                        //www.minix3.org/theses/
                                        veerman-dynamic-updates.
                                        pdf.

Vogt:2013:TEM

[VGBT13] Dirk Vogt, Cristiano Giuffrida, Herbert Bos, and Andrew S. Tanenbaum. Techniques for efficient in-memory checkpointing. In Christian Cachin and Robbert van Renesse, editors, Proceedings of HotDep ’13: the 9th Workshop on Hot Topics in Dependable Systems: November 13, 2013, Nemacolin Woodlands Resort, (Farmington, PA, USA), page ?? ACM Press, New York, NY 10036, USA, 2013. ISBN 1-4503-2457-6. LCCN QA76.9.F38. URL http:/

                                        /dl.acm.org/citation.cfm?
                                        id=2524224;   http://www.
                                        minix3.org/docs/workshop/
                                        hotdep-2013.pdf.

vanStaereling:2011:EMM

[vHvAvMT11] R. van Heuven van Staereling, Raja Appuswamy, David C. van Moolenbroek, and Andrew S. Tanenbaum. Efficient, modular metadata management with Loris. In IEEE, editor, Proceedings of the Sixth IEEE International Conference on Networking, Architecture, and Storage (NAS ’11), Dalian, Liaoning, China, 28–30 July 2011, pages 278– 287. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 0-7695-4509-2, 1-4577-1172-

9. LCCN TK5105.86. URL

http://ieeexplore.ieee.
org/servlet/opac?punumber=
6005103;   http://www.minix3.
org/docs/loris/NAS2011.
pdf.

Viken:1993:LRA

[Vik93] Jørn I. Viken. Lindex — en reimplementasjon av MINIX ved hjelp av Linda. (Norwegian) [Lindex: a reimplementation of MINIX with the help of Linda]. Hovedoppgave i informatikk (informatics thesis), Universitetet i Oslo, Oslo, Norway, 1993. 179 pp.

vanMoolenbroek:2007:MSM

[vM07] David van Moolenbroek. Multimedia support for MINIX

3. Master’s thesis in computer science, Department of Computer Science, Faculty of Sciences, Vrije Universiteit, Amsterdam, The Netherlands, September 2007. 187 pp.

URL http://www.minix3.
org/doc/moolenbroek_thesis.
pdf;   http://www.minix3.
org/theses/moolenbroek-
multimedia-support.pdf.

vanMoolenbroek:2011:IEE

[vMAT11] David C. van Moolenbroek, Raja Appuswamy, and Andrew S. Tanenbaum. Integrated end-to-end dependability in the Loris storage stack. In ????, editor, Proceedings of the Seventh Workshop on Hot Topics in System Dependability (HotDep’11), Hong Kong, China, June 2011, pages 165– 170. USENIX Association, Berkeley, CA, USA, 2011. ISBN ???? LCCN ????

URL http://www.minix3.
org/docs/loris/HotDep2011.
pdf.

vanMoolenbroek:2012:ISP

[vMAT12] David C. van Moolenbroek, Raja Appuswamy, and Andrew S. Tanenbaum. Integrated system and process crash recovery in the Loris storage stack. In IEEE, editor, NAS 2012: 2012 IEEE Seventh International Conference on Networking, Architecture, and Storage: proceedings: 28–30 June 2012, Xiamen, Fujian, China, pages 1–10. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-1889-2. LCCN TK5105.86.

URL http://www.minix3.
org/docs/loris/NAS2012.
pdf.

vanMoolenbroek:2013:BBB

[vMAT13a] David C. van Moolenbroek, Raja Appuswamy, and Andrew S. Tanenbaum. Battling bad bits with checksums in the Loris page cache. In IEEE, editor, LADC 2013: Sixth Latin-American Symposium on Dependable Computing: 2–5 April 2013, Rio de Janeiro, Brazil, pages 68–77. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN 0-7695-4962-4, 1-4673-5746-4. LCCN QA76.9.F3 L33 2013. URL

http://ieeexplore.ieee.
org/servlet/opac?punumber=
6542299;   http://www.minix3.
org/docs/conf/ladc-2013.

pdf. Best paper award.

vanMoolenbroek:2013:TBP

[vMAT13b] David C. van Moolenbroek, Raja Appuswamy, and Andrew S. Tanenbaum. Transactionbased process crash recovery of file system namespace modules. In IEEE, editor, Proceedings 2013 IEEE 19th Pacific Rim International Symposium on Dependable Computing: PRDC 2013: 2–4 December 2013: Vancouver, BC, Canada, pages 338–347. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN 0-7695-5130-0. LCCN QA76.9.F38. URL http://

www.minix3.org/docs/conf/
prdc-2013a.pdf.

vanMoolenbroek:2014:TFL

[vMAT14] David C. van Moolenbroek, Raja Appuswamy, and Andrew S. Tanenbaum. Towards a flexible, lightweight virtualization alternative. In ????, editor, SYSTOR ’14: proceedings of the 7th ACM International Systems and Storage Conference: June 10–12, 2014, Haifa, Israel, pages 1– 7. ACM Press, New York, NY 10036, USA, 2014. ISBN 1-4503-2920-9. LCCN ????

                                        URL http://dl.acm.org/
                                        citation.cfm?id=2611354;
                                         http://www.minix3.org/
                                        docs/conf/systor-2014.pdf.

Valero:1992:PCT

[VOJ^{+}92] M. Valero, E. Onate, M. Jane, J. L. Larriba, and B. Suarez, editors. Parallel Computing and Transputer Applications. CIMNE, Barcelona, Spain, 1992. ISBN 84-87867-13-8. LCCN ???? Two volumes.

Wainer:1995:IRT

[Wai95] Gabriel A. Wainer. Implementing real-time services in MINIX. Operating Systems Review, 29(3):75–84, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wei:1992:DSU

[Wei92] Yan Wei. Disk structures of Unix file systems. Mini-Micro Systems, 13(10):60–64, 1992. CODEN XWJXEH. ISSN 1000-1220.

Williams:1998:MLU

[Wil98] James D. Williams. A methodology for Linux as a user process based on Solaris Minix on the SPARC architecture.

Thesis (M.S.), New Mexico State University, as Cruces, NM 88003-8001, USA, 1998. xiii + 141 pp.

Winkler:1991:SPS

[Win91] Stanley Winkler, editor. Shortening the path from specification to prototype: the First International Workshop on Rapid System Prototyping, Research Triangle Park, North Carolina, USA, June 4–7, 1990. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-2175-3. LCCN QA76.9.C65 I577 1990. IEEE Cat. No.91TH0380-6.

Xu:1995:SIP

[Xu95] Li Xu. Study of an implementation of processes in MINIX. Thesis (m.s.), Mathematics and Computer Science, Central Missouri State University, Warrensburg, MO, USA, 1995. iv + 34 pp.

Yager:1990:AOSe

[Yag90] T. Yager. Alternative operating systems, part 5: Unix with a microscope: Minix, a low-cost Unix, runs on ordinary personal computers. Byte Magazine, 15(13):345– 346, December 1990. CODEN BYTEDJ. ISSN 0360-5280.

Yang:1995:SMO

[Yan95] Po-Tsun Yang. Study of MINIX operating system. Thesis (m.s.), Mathematics and Computer Science, Central Missouri State University, Warrensburg, MO 64093, USA, 1995. iv + 27 pp.

Yumerefendi:2005:RAD

[YC05] A. R. Yumerefendi and J. S. Chase. The role of accountability in dependable distributed systems. In ????, editor, Proceedings of the 1st Workshop on Hot Topics in System Dependability, page ?? USENIX Association, Berkeley, CA, USA, 2005. ISBN ???? LCCN ????